# A Survey on Federated Fine-Tuning of Large Language Models

### v2

@@ 5.2.1 General Evaluation Benchmarks
3) Instruction-following and generation under long input: LIFBENCH (Wu et al., 2024e), LongReason (Ling et al., 2025), BAMBOO (Dong et al., 2023b), ETHIC (Lee et al., 2024b), and LooGLE (Li et al., 2023e) evaluate multi-criteria instruction-following accuracy, response stability, and coherence under extended prompts. HELMET (Yen et al., 2024) further integrates summarization, retrieval, and reasoning in unified evaluation pipelines, supporting both automatic and human metrics. 4) Database-style and structured input reasoning: HoloBench (Maekawa et al., 2024) and LOFT (Lee et al., 2024a) benchmark the ability to perform complex reasoning over structured or database-like inputs, including table querying, execution accuracy, and factual consistency, especially in scenarios where retrieval-augmented methods are substituted by long-context modeling. 5) Scaling and generalization with longer context: Lv-Eval (Yuan et al., 2024b), ManyICLBench (Zou et al., 2024), ZeroSCROLLS (Shaham et al., 2023), and LongICLBench (Li et al., 2024f) examine the scalability of in-context learning and multi-task alignment as input length increases, making them valuable tools for analyzing the performance ceiling of long-context FedLLMs. 6) Cross-lingual long-context evaluation: LIBRA (Churin et al., 2024) tests instruction-following and long-form coherence in Russian, contributing to the evaluation of multilingual long-context capabilities. These long-context benchmarks are crucial for validating the scalability and persistence of FedLLMs, especially in environments requiring private document analysis, extended user sessions, or continual knowledge tracking. Table 10 presents a comprehensive summary of the general evaluation benchmarks discussed above.
Table 11: Overview of evaluation benchmarks across five specialized domains: finance, medicine, code, math, and law. The table summarizes representative benchmarks along with their primary evaluation objectives, target domains, and main evaluation criteria.
Benchmark Domain Evaluation Objective Main Evaluation Criteria
FinBen Finance Evaluate holistic financial capabilities across 24 tasks Automatic metrics, agent/RAG performance, and human expert assessment
PIXIU Finance Evaluate LLMs across multiple financial NLP tasks Sentiment accuracy, QA accuracy, stock prediction F1
FLUE Finance Evaluate diverse financial NLP competencies Accuracy, F1, nDCG
BBF-CFLEB Finance Evaluate LLMs on Chinese financial language understanding and generation across six task types Rouge, F1, and accuracy
CFinBench Finance Evaluate Chinese financial knowledge across subjects, certifications, practice, and legal compliance Accuracy across single-choice, multiple-choice, and judgment questions
SuperCLUEFin Finance Evaluate Chinese financial assistant capabilities Win-rate and multi-criteria performance
ICE-PIXIU Finance Evaluate bilingual (Chinese–English) financial reasoning and analysis capabilities Task-specific accuracy and bilingual win-rate
FLARE-ES Finance Evaluate bilingual Spanish–English financial reasoning Task-specific accuracy and cross-lingual transfer win-rate
FinanceBench Finance Evaluate financial open-book QA using real-world company-related questions Factual correctness and evidence alignment
FiNER-ORD Finance Evaluate financial NER capability in financial texts Entity F1
FinRED Finance Evaluate financial relation extraction performance on news and earnings transcripts F1, Entity F1
FinQA Finance Evaluate multi-step numerical reasoning over financial reports with structured evidence EM Accuracy
BizBench Finance Evaluate quantitative reasoning on realistic financial problems Numeric EM, code execution pass rate, QA accuracy
EconLogicQA Finance Evaluate economic sequential reasoning across multi-event scenarios Multiple-choice accuracy
FinEval Finance Evaluate Chinese financial domain knowledge and reasoning Multiple-choice accuracy and task-level weighted scores
CFBenchmark Finance Evaluate Chinese financial assistant capabilities Win-rate and task-specific metrics
BBT-Fin Finance Evaluate Chinese financial language understanding and generation Accuracy, F1, ROUGE
Hirano Finance Evaluate Japanese financial language understanding Multiple-choice accuracy and macro-F1 scores
MultiFin Finance Evaluate multilingual financial topic classification F1, Multi-class accuracy
DocFinQA Finance (long-context) Evaluate long-context financial reasoning over documents like financial reports EM and F1 for multi-step answer prediction, and reasoning accuracy
FinTextQA Finance (long-context) Evaluate long-form financial question answering with long textual context Answer accuracy, BLEU, and ROUGE
CBLUE Medical Evaluate Chinese biomedical language understanding across multiple clinical and QA tasks Accuracy, F1-score, and macro-average metrics across subtasks
PromptCBLUE Medical Evaluate LLMs on prompt-based generation across 16 Chinese medical NLP tasks Accuracy, BLEU, and ROUGE scores
CMB Medical Evaluate comprehensive Chinese medical knowledge via exam-style QA and clinical diagnosis tasks Accuracy, expert grading, and model-based evaluation
HuaTuo26M-test Medical Evaluate Chinese medical knowledge and QA ability using real-world clinical queries Accuracy and relevance
CMExam Medical Evaluate LLMs on Chinese medical licensing exam QA with fine-grained clinical annotations Accuracy, weighted F1 score, and expert-judged reasoning quality
MultiMedQA Medical Evaluate LLMs’ clinical knowledge via multiple-choice and open-ended medical QA tasks Expert-rated helpfulness, factuality, and safety
QiZhenGPT eval Medical Evaluate LLMs’ ability to identify drug indications from natural language prompts Expert-annotated correctness score
MedExQA Medical Evaluate LLMs’ medical knowledge and explanation generation across underrepresented specialties Explanation quality and expert-aligned LLM-judged relevance
JAMA and Medbullets Medical Evaluate LLMs’ ability to answer and explain challenging clinical questions Answer accuracy, explanation quality, and human-aligned reasoning assessment
MedXpertQA Medical Evaluate expert-level medical reasoning and multimodal clinical understanding across specialties Answer accuracy, image-text reasoning performance, and expert-aligned scoring
MedJourney Medical Evaluate LLM performance across full clinical patient journey stages and tasks Task-specific automatic metrics (e.g., accuracy) and human expert evaluations
MedAgentsBench Medical Evaluate complex multi-step clinical reasoning including diagnosis and treatment planning Multi-aspect evaluation including correctness, efficiency, and human expert ratings
LongHealth Medical (long-context) Evaluate question answering over long-form clinical documents Exact Match (EM), F1 score, and Long-Context QA accuracy
MedOdyssey Medical (long-context) Evaluate long-context understanding in the Medical domain Task-specific exact match, ROUGE, and human preference scoring
HumanEval Code Evaluate code generation, algorithmic reasoning, and language understanding with functional correctness Test-case execution accuracy (pass@k)
MBPP Code Evaluate basic Python code generation on crowdsourced programming tasks Functional correctness via pass@k using automated test cases
APPS Code Evaluate coding challenge competence through real-world programming problems Pass@k and exact match for functional correctness
DS-1000 Code Evaluate data science code generation across real-world queries from 7 Python libraries Functional correctness via automated test-based execution accuracy
CodeXGLUE Code Evaluate code understanding and generation across 9 tasks in 4 input-output types Automatic metrics including BLEU, EM, F1, Accuracy, and MAP
CruxEval Code Evaluate code reasoning, understanding, and execution pass@1 accuracy
ODEX Code Evaluate cross-lingual code generation from natural language queries in four languages Functional correctness via execution-based evaluation
MTPB Code Evaluate multi-turn program synthesis Functional correctness via pass@k on step-wise subprograms
ClassEval Code Evaluate class-level code generation from natural language descriptions in Python pass@1, class completeness, dependency consistency, and error analysis
BigCodeBench Code Evaluate LLMs’ ability to follow complex instructions and invoke diverse function calls Pass@k, test case execution accuracy, branch coverage
HumanEvalPack Code Evaluate multilingual code generation, correction, and comment synthesis across six programming languages Functional correctness via pass@k and task-specific auto metrics
BIRD Code Evaluate database-grounded text-to-SQL generation Execution accuracy and exact match
RepoQA Code (long-context) Evaluate long-context code understanding in real-world software repositories Exact match accuracy and retrieval-augmented correctness
LongCodeArena Code (long-context) Evaluate long-context code comprehension, generation, and editing across real-world tasks Task-specific accuracy, exact match, and human evaluation
GSM8K Math Assess grade school math reasoning Exact match accuracy
MATH Math Evaluate mathematical problem-solving ability on competition-level questions with step-by-step reasoning Final answer accuracy and step-wise derivation correctness
MathOdyssey Math Evaluate LLMs’ problem-solving capabilities on high school, university-level, and olympiad-level problems Answer accuracy and performance across varying difficulty levels
MathBench Math Evaluate theoretical understanding and practical application of mathematical knowledge across five levels Accuracy on theoretical and application problems
CHAMP Math Evaluate LLMs’ fine-grained mathematical reasoning with concept and hint annotations on competition-level problems Answer accuracy, reasoning path correctness
LILA Math Evaluate LLMs’ mathematical reasoning across diverse formats and topics Accuracy across 23 tasks
MiniF2F-v1 Math Evaluate formal mathematical reasoning at Olympiad level Proof accuracy on 488 problems
ProofNet Math Evaluate auto-formalization and formal proof generation in undergraduate-level mathematics Formalization accuracy and formal proof success rate in Lean 3
AlphaGeometry Math Evaluate neuro-symbolic reasoning on olympiad-level Euclidean geometry theorems Proof success rate, correctness, completeness, and human readability
MathVerse Math Evaluate MLLMs’ capability to interpret and reason over visual math diagrams Diagram-sensitive answer accuracy with fine-grained CoT reasoning score
We-Math Math Evaluate LMMs’ visual mathematical reasoning with emphasis on knowledge acquisition and generalization Four-dimensional diagnostic metrics
U-MATH Math Evaluate LLMs’ open-ended problem-solving skills across university-level math with visual components Solution correctness judged by LLMs with expert-verified F1 score
TabMWP Math Evaluate mathematical reasoning over textual and tabular data Accuracy on QA and multiple-choice questions
MathHay Math (long-context) Evaluate long-context mathematical reasoning with multi-step dependencies Accuracy, exact match, and reasoning chain consistency
LegalBench Law Evaluate legal reasoning across six types including rule application and interpretation Task-specific accuracy, rule-consistency, and LLM-as-a-judge ratings
LexGLUE Law Evaluate legal language understanding across classification and QA tasks Task-specific accuracy and F1 score (Auto)
LEXTREME Law Evaluate multilingual and multitask legal language understanding across 24 languages and 18 tasks Task-specific accuracy, macro-F1, and classification metrics (Auto)
LawBench Law Evaluate Chinese legal LLMs across retention, understanding, and application dimensions via 20 tasks Task-specific accuracy and F1
LAiW Law Evaluate Chinese legal LLMs across fundamental, basic, and advanced legal tasks Task-specific accuracy and F1 across 13 assignments
LexEval Law Evaluate LLMs’ Chinese legal understanding and reasoning using a taxonomy of legal cognitive abilities Task-specific accuracy
CitaLaw Law Evaluate LLMs’ ability to generate legally sound answers with appropriate citations using statutes and precedent cases Syllogism-based alignment score, citation accuracy and legal consistency
LegalAgentBench Law Evaluate LLM agents’ ability to solve complex real-world legal tasks Task success rate and intermediate progress rate
SCALE Law (long-context) Evaluate LLMs’ legal reasoning across languages, long documents, and multitask legal scenarios Accuracy, F1, and code-based assessment for long-context legal tasks
@@ 5.2.2 Financial Domain Evaluation Benchmarks
(i) Multi-task and agent-style financial evaluation. Several benchmarks offer broad-spectrum evaluation across multiple financial tasks, aligning well with FedLLM’s need to support diverse client use cases (e.g., auditing, compliance, trading). FinBen (Xie et al., 2024) evaluates holistic financial reasoning over 24 tasks using automatic metrics, retrieval-augmented generation accuracy, and expert human judgment. PIXIU (Xie et al., 2023), FLUE (Shah et al., 2022), and BBF-CFLEB (Lu et al., 2023a) benchmark LLMs on multi-task setups including sentiment classification, QA, event detection, and stock prediction. CFinBench (Nie et al., 2024) and SuperCLUEFin (Xu et al., 2024a) extend this paradigm to Chinese financial scenarios, assessing models across regulatory knowledge, certification preparation, and real-world task instructions using multi-type question formats. ICE-PIXIU (Xie et al., 2023) further supports bilingual Chinese–English evaluation, suitable for cross-regional FedLLM deployments. FLARE-ES (Zhang et al., 2024g) enables bilingual Spanish–English testing to evaluate cross-lingual transfer and domain-specific reasoning.
(ii) Task-specific capability evaluation. Fine-grained benchmarks evaluate specific financial NLP capabilities that are critical for real-world FedLLM deployment, including document question answering, information extraction, and numerical reasoning. FinanceBench (Islam et al., 2023) targets open-book question answering using real-world, company-related financial documents, with a focus on factual correctness and evidence alignment—a critical requirement for FedLLM deployed in compliance, auditing, and enterprise document analysis scenarios. FiNER-ORD (Shah et al., 2023) and FinRED (Sharma et al., 2022) evaluate financial Named Entity Recognition and relation extraction from news and earnings transcripts—key for localized data processing on devices with limited connectivity. FinQA (Chen et al., 2021b) and BizBench (Koncel-Kedziorski et al., 2023) benchmark multi-step numerical and quantitative reasoning, involving both tabular data and executable financial logic, useful in portfolio analysis, valuation, and budgeting applications. EconLogicQA (Quan & Liu, 2024) addresses sequential economic reasoning over multi-event scenarios. In the Chinese financial domain, FinEval (Zhang et al., 2023c), CFBenchmark (Lei et al., 2023), and BBT-Fin (Lu et al., 2023a) assess task-specific instruction following across topics such as taxation, accounting, and investment strategy. Hirano (Hirano, 2024) enables financial language understanding evaluation in Japanese, while MultiFin (Jørgensen et al., 2023) focuses on multilingual topic classification, useful in federated settings with cross-border clients or news sources.
@@ 5.2.3 Medical Domain Evaluation Benchmarks
(i) Medical knowledge and QA evaluation. This category focuses on assessing LLMs’ general and specialized medical knowledge through structured question-answering tasks. CBLUE (Zhang et al., 2021) and PromptCBLUE [^26] https://github.com/michael-wzhu/PromptCBLUE benchmark Chinese biomedical NLP tasks across information extraction, document classification, and QA, with the latter emphasizing prompt-based generation. CMB (Wang et al., 2023e), HuaTuo26M-test (Li et al., 2023d), and CMExam (Liu et al., 2023d) evaluate models on Chinese medical exam-style QA, clinical diagnosis tasks, and real-world query understanding. These benchmarks are well-suited for evaluating FedLLM tailored to localized clinical documentation and public health information. MultiMedQA (Singhal et al., 2023) serves as a high-quality English benchmark encompassing multiple-choice and open-ended medical QA, with expert-based evaluation of factuality, helpfulness, and safety—three pillars critical for deploying FedLLM in patient-facing applications. QiZhenGPT eval [^27] https://github.com/CMKRG/QiZhenGPT/tree/main/data/eval provides an annotation-based evaluation of drug indication extraction from natural language prompts, useful for drug interaction checking at the edge. MedExQA (Kim et al., 2024) expands QA evaluation to underrepresented medical specialties, while JAMA and Medbullets (Chen et al., 2025) challenge models with high-difficulty US medical exam-style questions and demand strong explanation quality.

### v3

@@ 5.2.1 General Evaluation Benchmarks
3) Instruction-following and generation under long input: LIFBENCH (Wu et al., 2024e), LongReason (Ling et al., 2025), BAMBOO (Dong et al., 2023b), ETHIC (Lee et al., 2024b), and LooGLE (Li et al., 2023e) evaluate multi-criteria instruction-following accuracy, response stability, and coherence under extended prompts. HELMET (Yen et al., 2024) further integrates summarization, retrieval, and reasoning in unified evaluation pipelines, supporting both automatic and human metrics. 4) Database-style and structured input reasoning: HoloBench (Maekawa et al., 2024) and LOFT (Lee et al., 2024a) benchmark the ability to perform complex reasoning over structured or database-like inputs, including table querying, execution accuracy, and factual consistency, especially in scenarios where retrieval-augmented methods are substituted by long-context modeling.
5) Scaling and generalization with longer context: Lv-Eval (Yuan et al., 2024b), ManyICLBench (Zou et al., 2024), ZeroSCROLLS (Shaham et al., 2023), and LongICLBench (Li et al., 2024f) examine the scalability of in-context learning and multi-task alignment as input length increases, making them valuable tools for analyzing the performance ceiling of long-context FedLLMs. 6) Cross-lingual long-context evaluation: LIBRA (Churin et al., 2024) tests instruction-following and long-form coherence in Russian, contributing to the evaluation of multilingual long-context capabilities. These long-context benchmarks are crucial for validating the scalability and persistence of FedLLMs, especially in environments requiring private document analysis, extended user sessions, or continual knowledge tracking. Table 11 and Table 12 summarize the general evaluation benchmarks discussed above.
@@ 5.2.2 Financial Domain Evaluation Benchmarks
(i) Multi-task and agent-style financial evaluation. Several benchmarks offer broad-spectrum evaluation across multiple financial tasks, aligning well with FedLLM’s need to support diverse client use cases (e.g., auditing, compliance, trading). FinBen (Xie et al., 2024) evaluates holistic financial reasoning over 24 tasks using automatic metrics, retrieval-augmented generation accuracy, and expert human judgment. PIXIU (Xie et al., 2023), FLUE (Shah et al., 2022), and BBT-CFLEB (Lu et al., 2023a) benchmark LLMs on multi-task setups including sentiment classification, QA, event detection, and stock prediction. CFinBench (Nie et al., 2024) and SuperCLUEFin (Xu et al., 2024a) extend this paradigm to Chinese financial scenarios, assessing models across regulatory knowledge, certification preparation, and real-world task instructions using multi-type question formats. ICE‑PIXIU (Xie et al., 2023) further supports bilingual Chinese–English evaluation, suitable for cross-regional FedLLM deployments. FLARE-ES (Zhang et al., 2024g) enables bilingual Spanish–English testing to evaluate cross-lingual transfer and domain-specific reasoning.
(ii) Task-specific capability evaluation. Fine-grained benchmarks evaluate specific financial NLP capabilities that are critical for real-world FedLLM deployment, including document question answering, information extraction, and numerical reasoning. FinanceBench (Islam et al., 2023) targets open-book question answering using real-world, company-related financial documents, with a focus on factual correctness and evidence alignment—a critical requirement for FedLLM deployed in compliance, auditing, and enterprise document analysis scenarios. FiNER‑ORD (Shah et al., 2023) and FinRED (Sharma et al., 2022) evaluate financial Named Entity Recognition and relation extraction from news and earnings transcripts—key for localized data processing on devices with limited connectivity. FinQA (Chen et al., 2021b) and BizBench (Koncel-Kedziorski et al., 2023) benchmark multi-step numerical and quantitative reasoning, involving both tabular data and executable financial logic, useful in portfolio analysis, valuation, and budgeting applications. EconLogicQA (Quan & Liu, 2024) addresses sequential economic reasoning over multi-event scenarios. In the Chinese financial domain, FinEval (Zhang et al., 2023c) and CFBenchmark (Lei et al., 2023) assess task-specific instruction following across topics such as taxation, accounting, and investment strategy. Hirano (Hirano, 2024) enables financial language understanding evaluation in Japanese, while MultiFin (Jørgensen et al., 2023) focuses on multilingual topic classification, useful in federated settings with cross-border clients.
@@
Table 13: Overview of evaluation benchmarks in the financial and medical domains.
Benchmark Domain Evaluation Objective Main Evaluation Criteria
FinBen Financial Holistic financial capabilities Auto metrics & Human eval
PIXIU Financial Financial NLP tasks Accuracy & F1
FLUE Financial Financial language understanding Accuracy, F1, nDCG, and MRR
BBT-CFLEB Financial Chinese financial understanding and generation ROUGE, F1, and Accuracy
CFinBench Financial Chinese financial knowledge Accuracy
SuperCLUEFin Financial Chinese financial assistant capabilities Win Rate & Accuracy
ICE‑PIXIU Financial Bilingual (Chinese–English) financial reasoning Accuracy, F1, and ROUGE
FLARE‑ES Financial Bilingual (Spanish–English) financial reasoning Accuracy, F1, and ROUGE
FinanceBench Financial Financial open-book QA (real-world) Factual Correctness
FiNER‑ORD Financial Financial NER F1, Precision, and Recall
FinRED Financial Financial RE (news, transcripts) F1, Precision, and Recall
FinQA Financial Numerical reasoning over financial reports Accuracy
BizBench Financial Quantitative reasoning on realistic problems Accuracy & F1
EconLogicQA Financial Economic sequential reasoning Accuracy
FinEval Financial Chinese financial knowledge and reasoning Accuracy
CFBenchmark Financial Chinese financial assistant capabilities LLM Judge Score & Accuracy
Hirano Financial Japanese financial understanding Accuracy
MultiFin Financial Multilingual financial topic classification F1 & Accuracy
DocFinQA Financial (LC) Long-context financial document reasoning Accuracy
FinTextQA Financial (LC) Long-form financial QA ROUGE & LLM Judge Score
CBLUE Medical Chinese biomedical understanding Accuracy & F1
PromptCBLUE Medical Chinese medical knowledge (prompt-based) Accuracy & F1
CMB Medical Chinese medical knowledge Accuracy & LLM Judge Score
HuaTuo26M Medical Chinese medical knowledge and QA ROUGE & and LLM Judge Score
CMExam Medical Chinese medical licensing exam Accuracy, BLEU, and ROUGE
MultiMedQA Medical Clinical knowledge and open-ended QA Accuracy & Human Evaluation
QiZhenGPT Medical Drug indication identification Accuracy
MedExQA Medical Medical knowledge and explanation generation Accuracy & LLM judge Score
JAMA and Medbullets Medical Challenging clinical QA Accuracy & Human Evaluation
MedXpertQA Medical Expert-level medical reasoning (multi-modal) Accuracy
MedJourney Medical Full clinical patient journey tasks Accuracy, ROUGE, BLEU, and F1
MedAgentsBench Medical Complex multi-step clinical reasoning Accuracy
LongHealth Medical (LC) QA over long-form clinical documents Accuracy
MedOdyssey Medical (LC) Long-context medical understanding Accuracy, Recall, and F1
@@ 5.2.3 Medical Domain Evaluation Benchmarks
(i) Medical knowledge and QA evaluation. This category focuses on assessing LLMs’ general and specialized medical knowledge through structured question-answering tasks. CBLUE (Zhang et al., 2021) and PromptCBLUE [^25] https://github.com/michael-wzhu/PromptCBLUE benchmark Chinese biomedical NLP tasks across information extraction, document classification, and QA, with the latter emphasizing prompt-based generation. CMB (Wang et al., 2023e), HuaTuo26M (Li et al., 2023d), and CMExam (Liu et al., 2023d) evaluate models on Chinese medical exam-style QA, clinical diagnosis tasks, and real-world query understanding. These benchmarks are well-suited for evaluating FedLLM tailored to localized clinical documentation and public health information. MultiMedQA (Singhal et al., 2023) serves as a high-quality English benchmark encompassing multiple-choice and open-ended medical QA, with expert-based evaluation of factuality, helpfulness, and safety—three pillars critical for deploying FedLLM in patient-facing applications. QiZhenGPT [^26] https://github.com/CMKRG/QiZhenGPT provides an annotation-based evaluation of drug indication extraction from natural language prompts, useful for drug interaction checking at the edge. MedExQA (Kim et al., 2024) expands QA evaluation to underrepresented medical specialties, while JAMA and Medbullets (Chen et al., 2025) challenge models with high-difficulty US medical exam-style questions and demand strong explanation quality.
@@ 5.2.4 Code Domain Evaluation Benchmarks
Table 14: Overview of evaluation benchmarks in the code, math, and law domains.
Benchmark Domain Evaluation Objective Main Evaluation Criteria
HumanEval Code Code generation and algorithmic reasoning Pass@k
MBPP Code Basic Python code generation Pass@k
APPS Code Coding challenge competence Pass@k
DS-1000 Code Data science code generation Accuracy & Pass@k
CodeXGLUE Code Code understanding and generation Accuracy, BLEU, and MRR
CruxEval Code Code reasoning, understanding, and execution Pass@k
ODEX Code Code generation and execution Pass@k
MTPB Code Multi-turn program synthesis Pass@k
ClassEval Code Class-level code generation Pass@k
BigCodeBench Code Complex and practical code generation Pass@k
HumanEvalPack Code Multilingual code tasks Pass@k
BIRD Code Database-grounded text-to-SQL Execution Accuracy & F1
RepoQA Code (LC) Long-context code understanding Pass Rate
LongCodeArena Code (LC) Long-context code tasks Pass@k, Accuracy, and EM
GSM8K Math Grade school math reasoning Accuracy
Competition MATH Math Competition math problem-solving Accuracy
MathOdyssey Math High-level math problem-solving (Olympiad) Accuracy
MathBench Math Comprehensive and hierarchical math evaluation Accuracy & Circular Evaluation
CHAMP Math Fine-grained math reasoning Accuracy
LILA Math Meta-benchmark for math reasoning Accuracy & Pass@k
MiniF2F Math Formal math reasoning (Olympiad) Proof Accuracy
ProofNet Math Auto-formalization and formal proof generation Accuracy & Success Rate
AlphaGeometry Math Neuro-symbolic reasoning (Euclidean geometry) Solve Rate
MathVerse Math Visual math reasoning Accuracy & CoT Score
We-Math Math Multi-step math reasoning Accuracy
U-MATH Math Open-ended visual math (university-level) LLM Judge Score
TabMWP Math Math reasoning over text and tabular data Accuracy
MathHay Math (LC) Long-context math reasoning (multi-step) Accuracy
LegalBench Law Legal reasoning across six tasks Accuracy, F1, and ROUGE
LexGLUE Law Legal language understanding Accuracy & F1
LEXTREME Law Multilingual/multitask legal understanding Accuracy & F1
LawBench Law Chinese legal reasoning Accuracy & F1
LAiW Law Chinese legal tasks Accuracy & F1
LexEval Law Chinese legal understanding and reasoning Accuracy
CitaLaw Law Legal answer generation MRR, Recall, and ROUGE
LegalAgentBench Law LLM agent legal task-solving Success Rate & LLM Judge Score
SCALE Law (LC) Multilingual/long-document legal reasoning ROUGE & METEOR